Circular No. 7523

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000el IN NGC 7290 T. Puckett, Mountain Town, GA; and D. George, Ottawa, ON, report

the discovery of an apparent supernova (mag 17.5) on an unfiltered CCD frame (limiting mag 20.3) taken with the Puckett Observatory 0.60-m automated supernova patrol telescope on Nov. 21.03 UT. The candidate is located at α = 22^{h}28^{m}26^{s}.62, δ = +17^{o}08′51′′.2 (equinox 2000.0), which is 1′′.5 east of the center of NGC 7290. The new object was also present on an unfiltered CCD frame taken by P. Boltwood, Stittsville, ON, with a 0.35-m reflector on Nov. 24.05. SN 2000el does not appear on unfiltered images taken by Puckett during June 1–Oct. 14, and it does not appear on Palomar Sky Survey images taken on 1995 July 30 (limiting mag ∼ 21.0) and on 1992 Sept. 1 (limiting mag ∼ 21.0).

KS 1947+300 A. Levine, Massachusetts Institute of Technology (MIT), and R. Cor-

bet, Goddard Space Flight Center (GSFC) and Universities Space Research Association, report on behalf of the RXTE ASM team at MIT and GSFC: “RXTE All-Sky Monitor observations of the transient x-ray source KS 1947+300 (Borozdin et al. 1990, Sov. Astron. Lett. 16, 345) show that the source is currently exhibiting an outburst that began around Oct. 23 and reached a maximum flux of ∼ 20 mCrab (1.5–12 keV) around Nov.

8. The source is at most weakly detected (flux < 6 mCrab) prior to the outburst. A search for periodic modulation in the pre-outburst light curve suggests a possible period of 41.7 ± 0.1 days with a semi-amplitude of ∼ 1 mCrab and an average maximum flux of ∼ 4 mCrab. The epoch of maximum flux is 1998 July 24.5 UT ± 4 days. These parameters would predict a maximum near Nov. 4. The optical counterpart of KS 1947+300 may be an early-type star (e.g., Goranskij et al. 1991, Sov. Astron. Lett. 17, 399). Further observations are encouraged.”

COMET 73P/SCHWASSMANN-WACHMANN 3 Total magnitude estimates (cf. IAUC 7518): Nov. 18.83 UT, 12.5 (K.

Kadota, Ageo, Saitama, Japan, 0.18-m reflector + CCD; 12′ tail in p.a. 299^{o}); 23.83, 10.4 (S. Yoshida, Ibaraki, Japan, 0.25-m reflector; visual).

2000 November 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green